OpenStudy (crisulcampo):

What is the volume of a spherical segment whose base has a height of 2 cm and the diameter of the sphere measures 6 cm... The answer says it's 28/3 *pi* cm^3 , but I didn't get that one... I need a guide, so I can even show my procedure

OpenStudy (crisulcampo):

I need to apply this formula= \[V=\frac{ 2 }{ 3 }\pi*r ^{2}h-\frac{ 1 }{ 3 }\pi*R ^{2}\left( r-h \right)\]

OpenStudy (crisulcampo):

@ShadowLegendX @TheSmartOne

OpenStudy (crisulcampo):

@Shad0w

OpenStudy (crisulcampo):

@skullpatrol

OpenStudy (3mar):

May I guide you?

OpenStudy (crisulcampo):

of coursee you can @3mar

OpenStudy (crisulcampo):

what is your advice?

OpenStudy (crisulcampo):

@zepdrix

OpenStudy (sshayer):



OpenStudy (sshayer):

you got it.

OpenStudy (crisulcampo):

take a look of my procedure...

OpenStudy (crisulcampo):

did you see?

OpenStudy (sshayer):

yes i have used the formula you gave. volume of sector of height 2=2/3 pi 3^2*2=36/3pi volume of cone =1/3 pi r^2 (R-h) =1/3 pi (sqrt 8)^2 (3-2)=8pi/3 then subtract. you need only volume ,no area.

OpenStudy (sshayer):

we need volume of cap only.

OpenStudy (sshayer):

radius of sphere 6/2=3 cm

OpenStudy (crisulcampo):

that's right! however I took h= 1 cm because exercise says that base is above centre "O" a distance of 2 cm, so "h" is the height that starts from base until the top of the sphere

OpenStudy (crisulcampo):

so, you say I should take as "h" value 2 cm instead of 1 cm right?

OpenStudy (crisulcampo):

thanks @sshayer

OpenStudy (3mar):

Sorry, I was not here! but I think @sshayer did it very well... thanks to him